please wrote out all steps NEATLY, thank you in advance
dy Use implicit differentiation to find dx 3y2 2x-5 2x + 5 dy dx =
Answer #1

Solution 3y2 = 2x-5 2x+5 y as y(x) Differentiate both sides of the equation with respect to x d (3y2) = (2+5) (3y?) = 3. (y)

2(2x+5)-2(2x-5) (2x+5) 20 20 (2x+5) 6y (y) = (2x+5) For convenience, write do (y) as y byy = (275 Divide both sides by 6y Oy